# Verbosity $\neq$ Veracity: Demystify Verbosity Compensation Behavior of Large Language Models

### v1

@@ Abstract
When unsure about an answer, humans often respond with more words than necessary, hoping that part of the response will be correct. We observe a similar behavior in large language models (LLMs), which we term “Verbosity Compensation” (VC). VC is harmful because it confuses the user understanding, leading to low efficiency, and influences the LLM services by increasing the latency and cost of generating useless tokens. In this paper, we present the first work that defines and analyzes Verbosity Compensation, explores its causes, and proposes a simple mitigating approach. We define Verbosity Compensation as the behavior of generating responses that can be compressed without information loss when prompted to write concisely. Our experiments, conducted on five datasets of knowledge and reasoning-based QA tasks with 14 newly developed LLMs, reveal three conclusions. 1) We reveal a pervasive presence of Verbosity Compensation across all models and all datasets. Notably, GPT-4 exhibits a VC frequency of 50.40%. 2) We reveal the large performance gap between verbose and concise responses, with a notable difference of 27.61% on the Qasper dataset. We also demonstrate that this difference does not naturally diminish as LLM capability increases. Both 1) and 2) highlight the urgent need to mitigate the frequency of VC behavior and disentangle verbosity with veracity. We propose a simple yet effective cascade algorithm that replaces the verbose responses with the other model-generated responses. The results show that our approach effectively alleviates the VC of the Mistral model from 63.81% to 16.16% on the Qasper dataset. 3) We also find that verbose responses exhibit higher uncertainty across all five datasets, suggesting a strong connection between verbosity and model uncertainty. Our dataset and code are available at https://github.com/psunlpgroup/VerbosityLLM.
@@ 1 Introduction
Throw everything at the wall and see what sticks.
— American Proverb
“Verbosity compensation” (VC) refers to a situation where someone is using excessively wordy language, attempting to make up for a lack of substance or clear communication by trying to appear more knowledgeable or persuasive, even if they are not adding significant meaning to the message. VC is widely studied in liguistic (Juola, 2008), education (Strunk Jr & White, 2007; Demir, 2019), and psychology (Brookshire & McNeil, 2014). Interestingly, we find that LLMs also exhibit Verbosity Compensation behaviors. Figure 1 shows a motivating example. In the first response, LLM generates a concise answer that is correct with low uncertainty. In the second and third responses, LLM tries to guess the correct answer by repeating the question, and producing ambiguity, leading to a verbose response with low performance and high uncertainty.
@@
In LLMs, the verbosity compensation behaviors are harmful. The reasons are twofold. First, for the users, the responses with verbosity compensation will lead to confusion and inefficiency (Fowler, 1927; Oppenheimer, 2006). Second, for the servers, the verbosity leads to unnecessary higher costs and higher latency. Besides, the performance of the model should not correlated to the length or format of responses because diverse users favor different lengths of responses. If verbose responses attain higher or lower performance than concise responses, verbosity compensation also leads to bias of users of different length preferences.
In this paper, we present the first study that defines, analyzes, and mitigates VC behavior. First, we define Verbosity Compensation (VC) as the behavior of generating responses that can be compressed without information loss when prompted to write concisely. To analyze the VC behavior systematically, we propose a benchmark by unifying four existing long-context question-answering datasets, and a reasoning-based language understanding dataset. We benchmark 14 newly proposed LLMs on five datasets and conduct a comprehensive analysis of the results. Next, we demystify the VC behavior in steps.
@@
We first obtain a fine-grained understanding of this behavior by manually inspecting the model-generated results and summarizing the pattern of VC by classifying them. Although we find that different models and datasets exhibit diverse distribution over types of VC behavior, we can categorize VC into five distinct types, including repeating questions, enumerating, ambiguity, verbose details, and verbose format. Next, we combine all these verbosity types and produce statistics of the frequency of LLM-generating VC responses, the result reveals a pervasive presence of verbosity compensation (VC) across all models and all datasets. Notably, GPT-4 exhibits a VC frequency of 50.40%. The high-frequency result amplifies the harmful influence of verbose responses. We further analyze the impact of VC on model performance. We define metrics that compute the difference between verbose and concise responses and apply these metrics to models. Our results indicate that verbose responses exhibit significantly different compared to concise ones, with a notable drop of 27.61% on the Qasper dataset, highlighting the urgent need to disentangle verbosity with veracity.
@@
Next, we investigate the underlying causes of VC. We use perplexity and Laplacian scores for open and closed-source models to measure the uncertainty of model responses. We find that verbose responses exhibit higher uncertainty across all five datasets, suggesting a strong connection between verbosity and model uncertainty. Last but not least, to mitigate VC in LLMs, we propose a simple yet effective cascade algorithm that replaces verbose responses with responses of larger LLMs. Experiments demonstrate the efficacy of our cascade algorithm through tests on three model combinations: Gemma to Gemini, Mistral to GPT-4, and Llama to Claude. The results show that our approach effectively alleviates the VC of the Mistral model from 63.81% to 16.16% on the Qasper dataset. We also modify the cascade algorithm for model routing and obtain significant improvement over the random selecting baseline.
@@
  • We define the verbosity compensation behavior of LLMs and propose a benchmark for evaluating the verbosity compensation phenomenon. We found LLMs suffer from significant verbosity.
  • We conduct a comprehensive analysis of the verbosity compensation and connect verbosity compensation behavior to 1) model performance and 2) model uncertainty.
@@ 3.1 Verbosity Compensation of LLMs
We first formalize the task as follows. A dataset $\mathcal{D}$ consists of multiple data samples where each consists of a source text $x$, a query $q$, and a ground truth $y$. A large language model $LLM(*)$ consumes the concatenation of $x,q$ and an instruction $I$ to produce the response $r$. For instruction $I$, we always ask LLM to generate as concisely as possible so that the model is instructed not to generate verbose responses. We call these tasks Concise Question Answering (CQA). Since the LLMs have maximum context window sizes $L_{c}$, we truncate the source to accommodate diverse context limits (details in  A.4).
We define a response $r$ to be verbose if and only if it can be further compressed with fewer tokens while keeping the same meaning. To detect whether the response $r$ is verbose or not, we also use a verbosity detector $V(x,y,r)$ which returns one if the response is verbose and vice versa. Based on the detector, verbosity compensation behavior for an LLM is defined as a triple $(x,y,r)$ where $V(x,y,r)=1$ to describe the behavior of generating a result that is more verbose than required. To quantify the frequency of verbosity compensation behavior, we define it as the ratio of verbose responses in each dataset $\sum_{(x,y)\in\mathcal{D}}V(x,y,r)/|\mathcal{D}|$.
@@ 3.2 Performance and Verbosity Compensation
$\Delta(\mathcal{D})=\frac{\sum_{(x,y)\in\mathcal{D}}V(x,y,r)\times\text{recall}(y,r)}{\sum_{(x,y)\in\mathcal{D}}V(x,y,r)}-\frac{\sum_{(x,y)\in\mathcal{D}}(1-V(x,y,r))\times\text{recall}(y,r)}{\sum_{(x,y)\in\mathcal{D}}(1-V(x,y,r))}$
@@ 4.1 Dataset Construction
For all these five datasets, to avoid the influence of gold length $y$ and easiness of measuring the verbosity, we only pick the samples where the gold answer contains less than $n$ words. We manually inspect the samples and find that when $n=4$, the answers are concise without descriptive context, and any answer longer than gold length is verbose. Thus, we set $n=4$ for all datasets.
For each dataset, if the number of sample is longer than 500, we randomly pick 500 samples from them. Otherwise, we keep the entire dataset. Finally, there are 449 samples in LongBench, 410 samples in NQ30 and 500 samples in Qasper, NarrativeQA, and MMLU.
@@ 4.2 Models
We use 14 LLMs in total across all experiments including both open-source and closed-source models in 6 families: GPT, Claude, Gemini, Llama, Gemma, Mistral. Details can be found in Appendix A.2. For each model, in addition to the prompt that introduce the task, we also ask them to “generate as concise as possible, use a single phrase if possible”. Since we constraint that all samples of the datasets contain only one to three tokens, we define $V(x,y,r)$ as $|r|>3$. Also, we select 30 samples from each dataset and conduct human inspections, finding that any response that is longer can be regarded as verbose.
@@ 5.1 Verbosity Compensation
After showing verbosity happens frequently in LLMs, we further conduct a human annotation to inspect verbose response patterns and classify them into five types. Specifically, we choose six combinations of model and dataset (Table 1) and pick out the samples with verbose responses that are not fully correct (recall $\neq 1$, $V(x,y,r)=1$). By checking all these samples, we classify verbosity compensation behavior into five types (Table 1): Ambiguity indicates not answering precisely; repeating question indicates repeating the tokens in the question or providing unrelated information; enumerating shows answering multiple answers in a row trying to cover the correct answer; verbose detail/format means generating more detailed explanations or format symbols. Then, we annotate the verbosity compensation behaviors and obtain statistics in diverse settings. As shown in Figure 3, the ratio distribution of five types of behavior varies across different models and datasets. Furthermore, the main type of gemini-1.5-flash is repeat questions on MMLU dataset (67.86%), and enumerating on Qasper dataset (47.62%). In contrast, llama-3-80b mainly produces verbose details on the Qasper dataset (32.87%). This shows that different datasets or models have a significantly different distribution of the main type of verbosity behavior.
@@ A.2 Details of Large Language Models
A.3 Experiment Details
During experiments, we use the default parameters of all 14 LLMs. We run gemma, llama, and mistral models from Huggingface [^4] https://huggingface.co/ on 8 A100 GPUs. For gpt, claude, and gemini models, we run with the official API from the official website. For all datasets, we use the same prompt show in Table 7.
@@ A.3 Experiment Details
A.4 Input Chunking Algorithm

### v2

@@ Abstract
Although Large Language Models (LLMs) have demonstrated their strong capabilities in various tasks, recent work has revealed LLMs also exhibit undesirable behaviors, such as hallucination and toxicity, limiting their reliability and broader adoption. In this paper, we discover an understudied type of undesirable behavior of LLMs, which we term Verbosity Compensation (VC) — similar to the hesitation behavior of humans under uncertainty — where they respond with excessive words such as repeating questions, introducing ambiguity, or providing excessive enumeration. We present the first work that defines and analyzes Verbosity Compensation, explores its causes, and proposes a simple mitigating approach. Our experiments, conducted on five datasets of knowledge and reasoning-based QA tasks with 14 newly developed LLMs, reveal three conclusions. 1) We reveal a pervasive presence of VC across all models and all datasets. Notably, GPT-4 exhibits a VC frequency of 50.40%. 2) We reveal the large performance gap between verbose and concise responses, with a notable difference of 27.61% on the Qasper dataset. We also demonstrate that this difference does not naturally diminish as LLM capability increases. Both 1) and 2) highlight the urgent need to mitigate the frequency of VC behavior and disentangle verbosity with veracity. We propose a simple yet effective cascade algorithm that replaces the verbose responses with the other model-generated responses. The results show that our approach effectively alleviates the VC of the Mistral model from 63.81% to 16.16% on the Qasper dataset. 3) We also find that verbose responses exhibit higher uncertainty across all five datasets, suggesting a strong connection between verbosity and model uncertainty. Our dataset and code are available at https://github.com/psunlpgroup/VerbosityLLM.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across various tasks, including reasoning (Hendrycks et al., 2021b; Cobbe et al., 2021), knowledge-based question answering (Dasigi et al., 2021; Yang et al., 2018), and planning (Valmeekam et al., 2023; Zheng et al., 2024). However, despite these impressive capabilities, numerous studies have highlighted that LLMs struggle with undesirable behaviors, such as hallucination (Huang et al., 2023), toxicity (Wen et al., 2023), and ethical bias (Tao et al., 2023), which can pose significant risks to users.
Recently, few studies have studied lengthy responses in LLMs, under chain-of-thought (Chiang & Lee, 2024; Nayab et al., 2024) and machine translation (Briakou et al., 2024) settings. However, these studies only analyze the lengthiness of the responses. Singhal et al. (2023) found the correlation between length and RLHF. However, the mechanism of decoding lengthy responses remains unknown.
In this paper, we discover an understudied undesirable behavior of LLMs. We term it Verbosity Compensation (VC) behavior. We define VC as generating responses that can be compressed without information loss when prompted to respond concisely. Instead of focusing merely on the lengthy issue, VC emphasizes the detailed behavior of generating compressible tokens with a low density of useful information. We also find VC is closely connected with the uncertainty of LLMs, demystifying the mechanism of the VC behavior, and improving the understanding of both VC and uncertainty.
@@
Interestingly, VC is similar to the hesitation behavior of humans under uncertainty (Juola, 2008; Brookshire & McNeil, 2014). Figure 1 shows a motivating example. In the first response, LLM generates a concise answer that is correct with low uncertainty. In the second and third responses, instead of generating an answer concisely, such as “16.5”, LLM repeats the question, and produces ambiguity, leading to a VC response with low performance and high uncertainty. VC is harmful and undesired for both users and servers. For the users, VC will lead to confusion and inefficiency (Fowler, 1927; Oppenheimer, 2006). When an LLM enumerates multiple answers, users are unclear which one is correct. Besides, VC also leads to bias among users of different length preferences if verbose answers attain higher/lower scores. For the servers, the verbosity leads to unnecessary higher costs and higher latency by generating useless tokens.
@@
To analyze the VC behavior systematically, we unify four long-context question-answering datasets and a reasoning-based language understanding dataset. Targeting observing indirect or low density of the information in the response rather than comparing the lengthiness of the result, we simply pick the samples with less than three tokens to easily judge VC behavior by counting the tokens in responses. We benchmark 14 newly proposed LLMs on proposed datasets. Although we find that different models and datasets exhibit diverse distribution, we can categorize VC into five distinct types, including repeating questions, enumerating, ambiguity, verbose details, and verbose format. The result reveals a pervasive presence of verbosity compensation (VC) across all models and all datasets. Notably, GPT-4 exhibits a VC frequency of 50.40%. Meanwhile, we found that verbose responses exhibit significantly different recall from concise ones, with a notable drop of 27.61% on the Qasper dataset, highlighting the urgent need to disentangle verbosity with veracity.
@@
Next, we measure the uncertainty of model responses using perplexity and Laplacian scores for open and closed-source models. We find that verbose responses exhibit higher uncertainty across all five datasets, suggesting a strong connection between verbosity and model uncertainty. Finally, we leverage the connection between performance and VC to develop a routing algorithm that obtains significant improvements over the random selecting baseline and uncertainty-based routing. To mitigate VC in LLMs, we propose a simple yet effective cascade algorithm that replaces verbose responses with responses of larger LLMs. Experiments demonstrate the efficacy of the proposed algorithm through tests on three model combinations: Gemma to Gemini, Mistral to GPT-4, and Llama to Claude. The results show that our approach effectively alleviates the VC of the Mistral model from 63.81% to 16.16% on the Qasper dataset. The insights above can inspire the development of practical applications and effective mitigation strategies. Future work can mitigate the uncertainty of the LLMs by alleviating VC behavior due to the proposed connections between them.
@@
  • We define VC and propose a comprehensive benchmark to evaluate 14 LLMs, revealing they suffer significantly from five types of VC.
  • We conduct a rigorous analysis and connect VC to 1) model performance and 2) model uncertainty, shedding light on future applications and mitigation.
@@ 3.1 Verbosity Compensation of LLMs
We first formalize the task as follows. A dataset $\mathcal{D}$ consists of multiple data samples where each consists of a source text $x$, a query $q$, and a ground truth $y$. A large language model $LLM(*)$ consumes the concatenation of $x,q$ and an instruction $I$ to produce the response $r$. We use $|r|$ to represent the tokens in $r$. For instruction $I$, we always ask LLM to generate as concisely as possible so that the model is instructed not to generate verbose responses. We call these tasks Concise Question Answering (CQA). Since the LLMs have maximum context window sizes $L_{c}$, we truncate the source to accommodate diverse context limits (details in  A.3).
We define a response $r$ to exhibit verbosity if and only if it can be further compressed with fewer tokens while keeping the same meaning. To detect VC, we define verbosity compensation detector $V(x,y,r)$ (abbreviated as the verbosity detector). Using this detector, VC behavior for an LLM is defined as a triple $(x,y,r)$ where $V(x,y,r)=1$, describing that the VC occurs in the response $r$. To quantify the frequency of VC, we define it as the ratio of VC responses in each dataset $\sum_{(x,y)\in\mathcal{D}}V(x,y,r)/|\mathcal{D}|$. It is worth noting that although $V(x,y,r)$ contains ground truth answer $y$, $y$ is usually used to provide additional information and the verbosity detectors can work reference-free.
@@ 3.2 Performance and Verbosity Compensation
$\Delta(\mathcal{D})=\frac{\sum_{(x,y)\in\mathcal{D}}(1-V(x,y,r))\times\text{recall}(y,r)}{\sum_{(x,y)\in\mathcal{D}}(1-V(x,y,r))}-\frac{\sum_{(x,y)\in\mathcal{D}}V(x,y,r)\times\text{recall}(y,r)}{\sum_{(x,y)\in\mathcal{D}}V(x,y,r)}$
@@ 4.1 Dataset Construction
For all these five datasets, to avoid the influence of gold length $y$ and the easiness of measuring the verbosity, we only pick the samples where the gold answer contains less than $n$ words. We manually inspect the samples and find that when $n=4$, the answers are concise without descriptive context, and any answer longer than gold length is verbose. Thus, we set $n=4$ for all datasets.
For each dataset, if the number of samples is longer than 500, we randomly pick 500 samples from them. Otherwise, we keep the entire dataset. Finally, there are 449 samples in LongBench, 410 samples in NQ30, and 500 samples in Qasper, NarrativeQA, and MMLU.
@@ 4.2 Models
We use 14 LLMs in total across all experiments including both open-source and closed-source models in 6 families: GPT, Claude, Gemini, Llama, Gemma, Mistral. Details can be found in Appendix A.2. For each model, in addition to the prompt that introduces the task, we also ask them to “generate as concisely as possible, use a single phrase if possible”. Since we constraint that all samples of the datasets contain only one to three tokens, we define $V(x,y,r)$ as $|r|>3$, meaning the number of the token in response $r$ is more than three. Also, we select 30 samples from each dataset and conduct human inspections, finding that any response that is longer can be regarded as verbose.
@@ 5.1 Verbosity Compensation
After showing verbosity happens frequently in LLMs, we further conduct a human annotation to inspect verbose response patterns and classify them into five types. Specifically, we choose six combinations of model and dataset (Table 1) and pick out the samples with verbose responses that are not fully correct (recall $\neq 1$, $V(x,y,r)=1$). By checking all these samples, we classify verbosity compensation behavior into five types (Table 1): Ambiguity indicates not answering precisely; repeating question indicates repeating the tokens in the question or providing unrelated information; enumerating shows answering multiple answers in a row trying to cover the correct answer; verbose detail/format means generating more detailed explanations or format symbols. Then, we annotate the verbosity compensation behaviors and obtain statistics in diverse settings. As shown in Figure 3, the ratio distribution of five types of behavior varies across different models and datasets. Furthermore, the main type of Gemini-1.5-flash is repeating questions on the MMLU dataset (67.86%), and enumerating on the Qasper dataset (47.62%). In contrast, llama-3-70b mainly produces verbose details on the Qasper dataset (32.87%). This shows that different datasets or models have a significantly different distribution of the main type of verbosity behavior.
@@ A.2 Details of Large Language Models
During experiments, we use the default parameters of all 14 LLMs. We run gemma, llama, and mistral models from Huggingface [^4] https://huggingface.co/ on 8 A100 GPUs. For gpt, claude, and gemini models, we run with the official API from the official website. For all datasets, we use the same prompt shown in Table 7. We design a reinforced prompt to ensure LLM understands concise responses are required. Thus, we reinforce the prompt by repetition, and explanation, especially for the weaker models, making a fairer comparison by avoiding failing to understand instructions. We evaluate the robustness of VC against diverse prompts in Apendix C.3.
@@
A.3 Input Chunking Algorithm
@@ B.4 Truncation Principle
Appendix C Supplementary Experiments
C.1 Comparison with Uncertainty-based Routing Algorithm
We further conduct an analysis to compare the performance of the proposed routing algorithm with the uncertainty-based routing algorithm in addition to the random baselines. For the uncertainty-based routing algorithm, we first use perplexity as the metric to rank the uncertainty of the responses generated by a small model. We select top Kuncertain samples and replace them with the responses generated by the larger model. We select K from a set of $\{0,10,20,\cdots,100\}$ and connect them to draw the curve in Figure 7. As can be seen, although the uncertainty-based routing algorithm can obtain a better performance than the random baseline, it is still worse than the proposed algorithm by comparing the AUC of the figure (Area Under the Curve), demonstrating the effectiveness of the proposed algorithm.
Figure 7: Routing performance of Mistral-7b to GPT-4o. X-axis (unit 10-3 dollars) is the average cost of running one sample. The Y-axis is the F-1 score averaged across the samples on one dataset. Routing performance (green line) is higher than the random baseline models (blue line) and uncertainty-based baseline (purple .
C.2 Verbosity Compensation in Trip Planning Dataset
To further demonstrate that VC generally occurs in diverse open-ended tasks with diverse response lengths, we run a trip planning dataset from the Natural-Plan benchmark Zheng et al. (2024) using two Llama-3 models and test VC frequency and performance gaps. The task is to find the itinerary regarding the order of visiting N cities. We randomly select 500 data points from the dataset to form our dataset. For the prompt design, we follow the zero-shot prompt in the original paper and add one sentence “Answer as concisely as possible, each step contains less than 10 words”. For the verbosity detector follows our CoT setting: $V(x,y,r)=\mathbbm{1}\left(\bigvee_{s\in S}|s|>10\right)$. The results are shown in Table 11. VC also occurs frequently in trip planning, demonstrating the general presence of VC in both short- and long-response open-ended tasks.
Table 11: VC frequency and performance gaps on trip planning dataset.
concise verbose $\Delta$ Avg. VC Freq.
llama-3-8b 15.18 3.62 11.56 9.22 51.49
llama-3-70b 21.81 4.87 16.94 19.63 12.87
C.3 Robustness of Verbosity Compensation against Prompt Choices
As shown in Table 7 We design a reinforced prompt to ensure LLM understands concise responses are required. Thus, we reinforce the prompt by repetition, explanation, etc., especially for the weaker models, making a fairer comparison by avoiding failing to understand instructions.
We further experiment with multiple possible prompts to show VC is not overfitting to certain prompt settings. We aim to show that as long as the model knows to generate as concise as possible, we can observe significant VC behaviors.
Table 12 shows the performance gap on MMLU and Qasper datasets using Llama-3-8b with different prompt designs. As can be seen, compared with the original prompt, the variation of the prompt can also observe a significant $\Delta$ over both datasets. This demonstrates the robustness of VC against the choice of prompts. It is worth noting that, “Answer as concise as possible” yields the highest scores on two datasets, as well as the highest $\Delta$, demonstrating a simpler prompt with less constraint might generate a larger performance gap between concise and verbose responses.
Table 12: Comparison between original and other variations of the prompts. VC consistently occurs, demonstrating the robustness of the VC against prompts.
MMLU Qasper
concise verbose $\Delta$ Avg. concise verbose $\Delta$ Avg.
Prompt in Table 7
58.4 44.82 13.57 55.6 58.99 54.6 4.39 55.98
Using a single phrase rather than a sentence. Please answer in 3 words.
55.13 43.43 11.71 52.70 54.22 48.11 6.11 51.30
Answer as concise as possible.
68.04 50.26 17.78 61.07 70.17 60.44 9.73 63.63
C.4 Evaluation of Verbosity and Performance on Same Test Instances
As shown in Table 2, and Table 3, the performance of concise and verbose samples is based on the split of the dataset. There is no overlap between the samples in the concise and verbose split. To prevent the influence of bias in different instances, we conduct an analysis that fixes the test instances and compares different models. Specifically, for each instance, we calculated the ratio of LLMs exhibiting VC behavior and reported the averaged ratio across datasets in Table 13. This approach also increases the robustness of our findings, as the support (number of samples) for each dataset is 14 times higher than when using a single model. As shown in the table, the performance $\delta$ is still pervasive for all five datasets. Specifically, on the Qasper dataset, the $\Delta$ reaches 16.22%
Table 13: Overall recall comparison between verbose and concise responses. Each dataset contains the prediction from all 14 LLMs.
concise verbose overall
Recall Support Recall Support $\Delta$ VC Freq. Avg. Recall
Qasper 61.85 2272 45.63 389 16.22 32.46 56.59
LongBench 50.31 1912 44.22 375 6.10 30.42 48.46
NarrativeQA 38.09 2540 31.67 355 6.42 36.29 35.76
MMLU 65.09 1694 51.47 475 13.62 24.20 61.79
NQ30 53.34 1516 44.89 362 8.45 26.41 51.10
C.5 Latency Comparison of CaSel Algorithm and Individule Models
We conduct an analysis to compare the useless token generated and the time cost of individual models and the CaSel algorithm on two datasets using Mistral-7b and GPT-4o. To assess the number of useless tokens generated, given a response $r$, we first define the useless tokens as the part with longer than 3 tokens in response $r$: $\sum_{i=1}^{N}\max(0,|r_{i}|-3),$ where $N$ is the number of samples in a dataset. As shown in Table 14, with our proposed cascade algorithm, the total inference time might be higher than using a small model (0.79 vs. 1.21 seconds per sample) and lower than using a large model (14.86 vs. 5.93 seconds per sample), but the number of useless tokens generated is much less. On the other hand, by using the proposed algorithm, the useless tokens generated decrease from 596/327 to 93, mitigating the VC rate from 41.40% to 21.00% on the NarrativeQA dataset, demonstrating that useless tokens greatly decrease by using the proposed algorithm.
Table 14: Comparison of the number of generated useless tokens and inference time. # Mistral/GPT indicates the number of useless tokens generated by Mistral-7b and GPT-4o on the dataset. # Total is the sum of # Mistal/GPT, showing the total number of useless tokens. Infer. Time is the running time of the algorithm per sample (Unit: second). CaSel (Mistral $\rightarrow$ GPT) generated the fewest number of useless tokens and maintained the lowest VC frequency. The inference time is higher than the small model but still lower than the larger model.
Qasper NarrativeQA
# Mistral # GPT # Total VC Freq. Infer. Time # Mistral # GPT # Total VC Freq. Infer. Time
Mistral-7b 663 N/A 663 63.81 0.80 596 N/A 596 41.40 1.22
GPT-4o N/A 207 207 31.79 1.27 N/A 327 327 50.40 14.86
Mistral $\rightarrow$ GPT 0 86 86 16.60 1.21 0 93 93 21.00 5.93
C.6 The Influence of the Digits in Responses
We analyze the performance and VC frequency of the samples with and without numbers using llama-3-8b on the Qasper and NarrativeQA dataset. The results are shown in Table 15. Although the model is easier to perform better on the sample without numbers, the VC frequency is relatively lower for the responses with digits. To understand the reason, we further inspect the Qasper dataset, we find that the samples with numbers are not as open-ended as the ones without numbers, meaning that the search space of the answers with numbers is smaller. This leads to a lower VC frequency and is easier to answer.
Table 15: Comparison between responses with digits and without digits. The responses with digits show lower verbosity compensation frequency.
Qasper NarrativeQA
concise verbose Avg. VC Freq. concise verbose Avg. VC Freq.
w/o digits 58.99 53.66 56.18 52.63 33.39 18.18 27.21 40.66
w/ digits 58.97 57.73 58.40 45.83 56.25 10.00 38.46 38.46
C.7 Response Length of Chain-of-Thought Experiments
Our evaluation is not limited to gold answers with less than 4 words. To demonstrate the generalization of the proposed VC behavior, we run the experiments on Chain-of-Though settings where the responses can contain more than 300 words. Table 16 shows the statistics of Chain-of-Thought experiments. The average response length can reach more than 50 words, and the VC behavior is still pervasive.
Table 16: Lengths of the generated responses under chain-of-thought setting. The maximum length of the generated results can reach more than 300 words demonstrating that VC occurs in long response settings.
MMLU Qasper
VC Freq. Min Len. Max Len. Avg Len. VC Freq. Min Len. Max Len. Avg Len.
gpt-3.5-turbo 51.49 3 90 26.24 37.62 4 81 23.38
gemma-2-9b 20.79 9 107 27.92 43.56 18 103 37.08
llama-3-8b 43.56 15 333 57.14 44.15 20 185 50.15